# A Sanity Check for AI-generated Image Detection

### v2

@@ 3.2.2 Dataset Curation
To ensure the collection of high-quality images, we implement a comprehensive pipeline for image cleaning: (i) we discard images with resolution lower than $448\times 448$, as higher-resolution images generally provide better assessments of the robustness of existing models; (ii) due to the potential presence of violent and inappropriate content, we utilize SD’s safety checker model (saf, 2022) to filter out NSFW images; (iii) to avoid duplicated images, we compare their hash values to filter out duplicated images. In addition to this general cleaning pipeline, we introduce CLIP (Radford et al., 2021) to further filter out images with low image-text similarity. Specifically, for fake images, the online website provides prompts used to generate these images, and we calculate similarity using their corresponding prompts. For real images, we used the mean of the 80 prompt templates (e.g., a photo of {category} and a photo of the {category}) evaluated in CLIP’s ImageNet zero-shot as the text embedding.
3.2.3 Dataset Annotation
At this stage, we establish an annotation platform and recruit 20 human workers to manually label each of the AI-generated images for their category and realism. For categorization, annotators are instructed to assign each image to one of four major categories: human, animal, object, and scene. Regarding realism assessment, workers are tasked with labeling the images as either Real or AI-generated, based on the criterion of “whether this image could be taken with a camera”. It’s important to note that as the annotators are not informed whether the images are generated by AI algorithms beforehand. Each image was assessed independently by two annotators, and those have been misclassified as real by both annotators can thus be considered to pass the “perception turing test” and labeled as “highly realistic”. Subsequently, we retain only those images judged as “highly realistic”. Similarly, for real images, we follow the same procedure, retaining only those belonging to the four predefined categories, as we have done for AI-generated images.
3.2.4 Dataset Comparison
Our objective is to construct a sophisticated and exhaustive test dataset for AI-generated image detection. In Table 2, we conduct a comparative analysis between our Chameleon dataset and existing datasets. Our dataset is characterized by three pivotal features: Magnitude. Comprising about 26,000 test images, the Chameleon dataset represents the most extensive collection available, enhancing its robustness. Variety. Our dataset incorporates images from a vast array of real-world scenarios, surpassing the limited categorical focus of other datasets. Resolution. With resolutions spanning from 720P to 4K, With image resolutions ranging from 720P to 4K, artifacts demand more nuanced analysis, thus presenting additional challenges to the model due to the need for fine-grained discernment. In summary, our dataset offers a more demanding and pragmatically relevant benchmark for the advancement of AI-generated image detection methodologies.
Table 2: Comparison of AI-generated image detection testset. Our Chameleon dataset is the first to encompass real-life scenarios for detector evaluation. Compared to AIGCDetectBenchmark (Zhong et al., 2023), Chameleon offers greater magnitude and superior quality, rendering it more realistic in evaluation.
@@ 3.2.4 Dataset Comparison
Variety LSUN LUSN ImageNet ImageNet CelebA COCO LSUN FFHQ ImageNet ImageNet ImageNet ImageNet ImageNet ImageNet ImageNet ImageNet Real-life
@@ 4 Methodology
In this section, we present AIDE (AI-generated Image DEtector with Hybrid Features), consisting of a module to compute patchwise low-level statistics of texture or smooth patches, a high-level semantic embedding module, and a discriminator to classify the image as being generated or photographed. The overview of our AIDE model is illustrated in Fig. 2.
@@ 4.1 Patchwise Feature Extraction
We leverage insights from the disparities in low-level patch statistics between AI-generated images and real-world scenes. Models like generative adversarial networks or diffusion models often yield images with certain artifacts, such as excessive smoothness or anti-aliasing effects. To capture such discrepancy, we adopt a Discrete Cosine Transform (DCT) score module to identify patches with the highest and lowest frequency. By focusing on these extreme patches, we aim to highlight the distinctive characteristics of AI-generated images, thus facilitating the discriminative power of our detection system.
@@
To acquire the highest and lowest image patches, we use the complexity of the frequency components as an indicator. From this, we design a simple yet effective scoring mechanism. Specifically, we design $K$ different band-pass filters:
@@
where $F_{k,ij}$ is the weight at the $(i,j)$ position of the $k$-th filter. Next, for $m$-th patch $x_{m}^{\text{dct}}\in\mathbb{R}^{N\times N\times 3}$, we apply the filters $F_{k,ij}\in\mathbb{R}^{N\times N\times 3}$ to multiply the logarithm of the absolute DCT coefficients $x_{m}^{\text{dct}}\in\mathbb{R}^{N\times N\times 3}$ and sum all the positions to obtain the grade of the patch $G^{m}$. We formulated it as
@@
Patchwise Feature Encoder. Next, firstly, these patches are resized to a size of $256\times 256$. Secondly, they are input into the SRM (Fridrich & Kodovsky, 2012) to extract their noise pattern. Lastly, these features are input into two ResNet-50 (He et al., 2016) networks ($f_{1}(\cdot)$ and $f_{2}(\cdot)$) to obtain the final feature map $F_{\text{max}}=\{f_{1}(X_{\text{max}_{1}}),f_{1}(X_{\text{max}_{2}}),...,f_{1}(X_{\text{max}_{k}})\}$, $F_{\text{min}}=\{f_{2}(X_{\text{min}_{1}}),f_{2}(X_{\text{min}_{2}}),...,f_{2}(X_{\text{min}_{k}})\}$. We acquire the highest frequency embedding and lowest frequency embedding on the mean-pooled feature:
$F_{\text{\text{max}}}=\text{Mean}(\text{AveragePool}(F_{\text{max}})),~{}~{}~{}~{}F_{\text{\text{min}}}=\text{Mean}(\text{AveragePool}(F_{\text{min}})).$ (5)
@@ 4.3 Discriminator
To distinguish between AI-generated images and real images, we utilize a mixture-expert-model for the final discrimination. At low-level, we take the average of the highest frequency featured:
@@
$y=f([\text{avgpool}(F_{\text{mean}};F_{\text{s}}]),$ (8)
@@ 5.1 Experimental Details
Detectors. We evaluate 9 off-the-shelf detectors including CNNSpot (Wang et al., 2020), FreDect (Frank et al., 2020), Fusing (Ju et al., 2022), LNP (Liu et al., 2022a), LGrad (Tan et al., 2023), UnivFD (Ojha et al., 2023), DIRE (Wang et al., 2023), PatchCraft (Zhong et al., 2023) and NPR (Tan et al., 2024) for comparison.
@@
Implementation Details. AIDE includes two key modules: Patchwise Feature Extraction (PFE) and Semantic Feature Embedding (SFE). For PFE channel, we first patchify each image into patches and the patch size is set to be $N=32$ pixels. Then these patches are sorted using our DCT Scouring module with $K=6$ different band-pass filters in the frequency domain. Subsequently, we select two highest-frequency and two lowest-frequency patches using the calculated DCT scores. These selected patches are then resized to $256\times 256$ and extracted their noise pattern using SRM (Fridrich & Kodovsky, 2012). For SFE channel, we use the pre-trained OpenCLIP (Ilharco et al., 2021) to extract semantic features. We adopt data augmentations including random JPEG compression (QF $\sim\text{Uniform}(30,100)$) and random Gaussian blur ($\sigma\sim\text{Uniform}(0.1,3.0)$) to improve the robustness of detectors. Each augmentation is conducted with 10% probability. During the training phase, we use AdamW optimizer with the learning rate of $1\times 10^{-4}$ in $\mathcal{B}_{1}$ and $5\times 10^{-4}$ in $\mathcal{B}_{2}$, respectively. The batch size is set to $32$ and the model is trained on 8 NVIDIA A100 GPUs for only 5 epochs. Our method trains very quickly, only 2 hours are sufficient.
Metrics. In accordance with existing AI-generated detection arpproaches (Wang et al., 2020; 2019; Zhou et al., 2018), we report both classification accuracy (Acc) and average precision (AP) in our experiments. All results are averaged over both real and AI-generated images unless otherwise specified. We primarily report Acc for evaluation and comparison in the main paper, and AP results are presented in the Appendix.
5.2 Comparison to State-of-the-art Models
Figure 3: Performance of SOTA method, PatchCraft, under $\mathcal{B}_{1}$ (left), $\mathcal{B}_{2}$ (right), and our Chameleon testset. The boundary line for Acc = 50% is marked with a dashed line.
AIDE demonstrates a significant advancement over the current state-of-the-art (SOTA) approach, PatchCraft, achieving an average accuracy increase of 3.5%. UnivFD utilizes CLIP semantic features for detecting AI-generated images, proving effective for GAN-generated images. However, it shows pronounced performance degradation with diffusion model (DM)-generated images. This suggests that as generation quality improves, diffusion models produce images with fewer semantic artifacts, as depicted in Fig. 1 (a). Our approach, which integrates semantic, low-frequency, and high-frequency information at the feature level, enhances detection performance, yielding a 5.2% increase for GAN-based images and a 1.7% increase for DM-based images compared to the SOTA method.
Table 3: Comparison on the AIGCDetectBenchmark (Zhong et al., 2023) benchmark. Accuracy (%) of different detectors (rows) in detecting real and fake images from different generators (columns). DIRE-D indicates this result comes from DIRE detector trained over fake images generated by ADM following its official setup (Wang et al., 2023). DIRE-G indicates this baseline is trained on the same ProGAN training data as others. The best result and the second-best result are marked in bold and underline, respectively.
@@ 5.2 Comparison to State-of-the-art Models
Table 4: Comparison on the GenImage (Zhu et al., 2024) benchmark. Accuracy (%) of different detectors (rows) in detecting real and fake images from different generators (columns). These methods are trained on real images from ImageNet and fake images generated by SD v1.4 and evaluated over eight generators. The best result and the second-best result are marked in bold and underline, respectively.
@@
Table 5: Comparisons on the Chameleon benchmark. Accuracy (%) of different detectors (rows) in detecting real and fake images of Chameleon. For each training dataset, the first row indicates the Acc evaluated on the Chameleon testset, and the second row gives “fake image Acc / real image Acc” for detailed analysis.
@@
ProGAN 56.94 55.62 56.98 58.94 57.11 57.22 58.19 53.76 57.29 56.45
0.08/99.67 13.72/87.12 0.01/99.79 4.76/99.66 0.09/99.97 3.18/97.83 3.25/99.48 1.78/92.82 2.20/98.70 0.63/98.46
SD v1.4 60.11 56.86 57.07 60.95 55.63 55.62 59.71 56.32 58.13 61.10
8.86/98.63 1.37/98.57 0.00/99.96 17.65/93.50 0.57/97.01 74.97/41.09 11.86/95.67 3.07/96.35 2.43/100.00 16.82/94.38
All GenImage 60.89 57.22 57.09 59.81 58.52 60.42 57.83 55.70 57.81 63.89
9.86/99.25 0.89/99.55 0.02/99.98 8.23/98.58 7.72/96.70 85.52/41.56 2.09/99.73 1.39/96.52 1.68/100.00 22.40/95.06
On Benchmark Chameleon: As highlighted in Sec. 1, we contend that success on existing public benchmarks may not accurately reflect real-world scenarios or the advancement in AI-generated image detection, given that test sets are typically randomly sampled from generative models without ”Turing Test”. To address potential biases related to training setups—such as generator types and image categories—we evaluate the performance of existing detectors under diverse training conditions. Despite their high performance on existing benchmarks, as depicted in Fig. 3, the state-of-the-art detector, PatchCraft, experiences substantial performance declines. Additionally, Table 5 reveals significant performance decreases across all methods, with most struggling to surpass an average accuracy close to random guessing (about 50%), indicating a failure in these contexts.
While our method achieves state-of-the-art results on available datasets, its performance on Chameleon remains lacking. This underscores that our dataset, Chameleon, which challenges human perception, represents a critical issue requiring attention in this field.
Table 6: Robustness on JPEG Compression and Gaussian Blur of AIDE. The classification accuracy (%) averaged over 16 test sets in $\mathcal{B}_{1}$ with specific perturbation.
Method JPEG Compression Gaussian Blur
QF=95 QF=90 $\sigma=1.0$ $\sigma=2.0$
CNNSpot 64.03 62.26 68.39 67.26
FreDect 66.95 67.45 65.75 66.48
Fusing 62.43 61.39 68.09 66.69
GramNet 65.47 64.94 68.63 68.09
LNP 53.58 54.09 67.91 66.42
LGrad 51.55 51.39 71.73 69.12
DIRE-G 66.49 66.12 64.00 63.09
UnivFD 74.10 74.02 70.31 68.29
PatchCraft 72.48 71.41 75.99 74.90
AIDE 75.54 74.21 81.88 80.35
5.3 Robustness to Unseen Perturbations
In real-world scenarios, images often encounter unseen perturbations during transmission and interaction, complicating the detection of AI-generated images. Here, we assess the performance of various methods in handling potential perturbations, such as JJPEG compression (Quality Factor (QF) = 95, 90) and Gaussian blur ($\sigma$ = 1.0, 2.0). As illustrated in Table 6, all methods exhibit a decline in performance due to disruptions in the pixel distribution. These disruptions diminish the discriminative artifacts left by generative models, complicating the differentiation between real and AI-generated images. Consequently, the robustness of these detectors in identifying AI-generated images is significantly compromised. Despite these challenging conditions, our method consistently outperforms others, maintaining a relatively higher accuracy in detecting AI-generated images. This superior performance is attributed to our model’s ability to effectively capture and leverage multi-perspective features, semantics and noise, even when the pixel distribution is distorted.
5.4 Ablation Studies
Figure 4: Visualization of the effectiveness of PFE and SFE Modules. The absence of our semantic feature extraction module results in numerous AI-generated images exhibiting pronounced semantic errors that are incorrectly classified as real. Similarity, the second row demonstrates that when the patchwise feature extraction module is omitted, many AI-generated images, despite lacking semantic errors, contain subtle underlying noise that also leads to their misclassification as true.
Table 7: Ablation studies on Patchwise Feature Extraction and Semantic Feature extraction of AIDE.

### v3

@@ 3.2.2 Dataset Curation
To ensure the collection of high-quality images, we implement a comprehensive pipeline for image cleaning: (i) we discard images with resolution lower than $448\times 448$, as higher-resolution images generally provide better assessments of the robustness of existing models; (ii) due to the potential presence of violent and inappropriate content, we utilize SD’s safety checker model (saf, 2022) to filter out NSFW images; (iii) to avoid duplicated images, we compare their hash values to filter out duplicated images. In addition to this general cleaning pipeline, we introduce CLIP (Radford et al., 2021) to further filter out images with low image-text similarity. Specifically, for AI-generated images, the online website provides prompts used to generate these images, and we calculate similarity using their corresponding prompts. For real images, we used the mean of the 80 prompt templates (e.g., a photo of {category} and a photo of the {category}) evaluated in CLIP’s ImageNet zero-shot as the text embedding.
Table 2: Comparison of AI-generated image detection testset. Our Chameleon dataset is the first to encompass real-life scenarios for evaluation. Compared to AIGCDetectBenchmark (Zhong et al., 2023), Chameleon offers greater magnitude and superior quality, rendering it more realistic in evaluation. IN represents ImageNet.
@@
Variety LSUN LUSN IN IN CelebA COCO LSUN FFHQ IN IN IN IN IN IN IN IN Real-life
3.2.3 Dataset Annotation
At this stage, we establish an annotation platform and recruit 20 human workers to manually label each of the AI-generated images for their category and realism. For categorization, annotators are instructed to assign each image to one of four major categories: human, animal, object, and scene. Regarding realism assessment, workers are tasked with labeling the images as either Real or AI-generated, based on the criterion of whether this image could be taken with a camera. Note that, the annotators are not informed whether the images are generated by AI algorithms beforehand. Each image was assessed independently by two annotators, and those have been misclassified as real by both annotators can thus be considered to pass the “perception turing test” and labeled as highly realistic. Subsequently, we retain only those images judged as highly realistic. Similarly, for real images, we follow the same procedure, retaining only those belonging to the four predefined categories, as we have done for AI-generated images.
3.2.4 Dataset Comparison
Our objective is to construct a sophisticated and exhaustive test set that serves as a valuable extension to the current evaluation benchmarks for AI-generated image detection. In Table 2, we conduct a comparative analysis between our Chameleon dataset and existing test sets. Our dataset is characterized by three pivotal features: (i) Magnitude. Comprised of approximately 26,000 test images, the Chameleon dataset represents the most extensive collection available, surpassing any existing test set and enhancing its robustness. (ii) Variety. Our dataset incorporates images from a vast array of real-world scenarios, surpassing the limited categorical focus of the other datasets. (iii) Resolution. With resolutions spanning from 720P to 4K, With image resolutions ranging from 720P to 4K, artifacts demand more nuanced analysis, thus presenting additional challenges to the model due to the need for fine-grained discernment. In summary, our dataset offers a more demanding and pragmatically relevant benchmark for the advancement of AI-generated image detection methodologies.
@@ 4 Methodology
In this section, we present AIDE (AI-generated Image DEtector with Hybrid Features), consisting of a module to compute low-level statistics for texture or smooth patches, a high-level semantic embedding module, and a discriminator to classify the image as being generated or photographed. The overview of our AIDE model is illustrated in Fig. 2.
@@ 4.1 Patchwise Feature Extraction
Here, our design leverages the disparities in low-level patch statistics between AI-generated images and real-world scenes. Models like generative adversarial networks or diffusion models often yield images with certain artifacts, such as excessive smoothness or anti-aliasing effects. To capture such discrepancy, we adopt a Discrete Cosine Transform (DCT) score module to identify patches with the highest and lowest frequency. By focusing on these extreme patches, we aim to highlight the distinctive characteristics of AI-generated images, thus facilitating the discriminative power of our detection system.
@@
To acquire the highest and lowest image patches, we use the complexity of the frequency components as an indicator. From this, we design a simple yet effective scoring mechanism, using $K$ different band-pass filters:
@@
where $F_{k,ij}$ is the weight at the $(i,j)$ position of the $k$-th filter. Next, for the $m$-th patch ($x_{m}^{\text{dct}}\in\mathbb{R}^{N\times N\times 3}$), we apply the filters $F_{k,ij}\in\mathbb{R}^{N\times N\times 3}$ to multiply the logarithm of the absolute DCT coefficients ($x_{m}^{\text{dct}}\in\mathbb{R}^{N\times N\times 3}$) and sum up all the positions to obtain the grade of the patch $G^{m}$. We formulated it as
@@
Patchwise Feature Encoder. These patches are resized to size of $256\times 256$ pixels, and input into the SRM (Fridrich & Kodovsky, 2012) to extract their noise pattern. Lastly, these features are further input into two ResNet-50 (He et al., 2016) networks ($f_{1}(\cdot)$ and $f_{2}(\cdot)$) to obtain the final feature map $F_{\text{max}}=\{f_{1}(X_{\text{max}_{1}}),f_{1}(X_{\text{max}_{2}}),...,f_{1}(X_{\text{max}_{k}})\}$, $F_{\text{min}}=\{f_{2}(X_{\text{min}_{1}}),f_{2}(X_{\text{min}_{2}}),...,f_{2}(X_{\text{min}_{k}})\}$. We acquire the highest and lowest frequency embedding on the mean-pooled feature:
$F_{\text{\text{max}}}=\text{Mean}(\text{AveragePool}(F_{\text{max}})),~{}~{}~{}~{}F_{\text{\text{min}}}=\text{Mean}(\text{AveragePool}(F_{\text{min}}))$ (5)
@@ 4.3 Discriminator
To distinguish between AI-generated images and real images, we utilize a mixture-expert model for the final discrimination. At low-level, we take the average of the highest frequency featured:
@@
$y=f([\text{avgpool}(F_{\text{mean}};F_{\text{s}}])$ (8)
@@ 5.1 Experimental Details
Baseline Detectors. We evaluate 9 off-the-shelf detectors including CNNSpot (Wang et al., 2020), FreDect (Frank et al., 2020), Fusing (Ju et al., 2022), LNP (Liu et al., 2022a), LGrad (Tan et al., 2023), UnivFD (Ojha et al., 2023), DIRE (Wang et al., 2023a), PatchCraft (Zhong et al., 2023) and NPR (Tan et al., 2024) for comparison.
@@
Implementation Details. AIDE includes two key modules: Patchwise Feature Extraction (PFE) and Semantic Feature Embedding (SFE). For PFE channel, we first patchify each image into patches and the patch size is set to be $N=32$ pixels. Then these patches are sorted using our DCT Scoring module with $K=6$ different band-pass filters in the frequency domain. Subsequently, we select two highest-frequency and two lowest-frequency patches using the calculated DCT scores. These selected patches are then resized to $256\times 256$ and extracted their noise pattern using SRM (Fridrich & Kodovsky, 2012). For SFE channel, we use the pre-trained OpenCLIP (Ilharco et al., 2021) to extract semantic features. We adopt data augmentations including random JPEG compression (QF $\sim\text{Uniform}(30,100)$) and random Gaussian blur ($\sigma\sim\text{Uniform}(0.1,3.0)$) to improve the robustness of detectors. Each augmentation is conducted with 10% probability. During the training phase, we use AdamW optimizer with the learning rate of $1\times 10^{-4}$ in $\mathcal{B}_{1}$ and $5\times 10^{-4}$ in $\mathcal{B}_{2}$, respectively. The batch size is set to $32$ and the model is trained on 8 NVIDIA A100 GPUs for only 5 epochs. Our method trains very quickly, only 2 hours are sufficient.
Metrics. In accordance with existing AI-generated detection approaches Wang et al. (2020; 2019); Zhou et al. (2018), we report both classification accuracy (Acc) and average precision (AP) in our experiments. All results are averaged over both real and AI-generated images unless otherwise specified. We primarily report Acc for evaluation and comparison in the main paper, and AP results are presented in the Appendix.
Table 3: Comparison on the AIGCDetectBenchmark (Zhong et al., 2023). Accuracy (%) of different detectors (rows) in detecting real and fake images from different generators (columns). DIRE-D indicates this result comes from DIRE detector trained over fake images generated by ADM following its official setup (Wang et al., 2023a). DIRE-G indicates this baseline is trained on the same ProGAN training data as others. The best result and the second-best result are marked in bold and underline, respectively.
@@
Table 4: Comparison on the GenImage (Zhu et al., 2024). Accuracy (%) of different detectors (rows) in detecting real and fake images from different generators (columns). These methods are trained on real images from ImageNet and fake images generated by SD v1.4 and evaluated over eight generators. The best result and the second-best result are marked in bold and underline, respectively.
@@
5.2 Comparison to State-of-the-art Models
Table 5: Comparison on the Chameleon. Accuracy (%) of different detectors (rows) in detecting real and fake images of Chameleon. For each training dataset, the first row indicates the Acc evaluated on the Chameleon testset, and the second row gives the Acc for “fake image / real image” for detailed analysis.
@@ 5.2 Comparison to State-of-the-art Models
ProGAN 56.94 55.62 56.98 58.94 57.11 57.22 58.19 53.76 57.29 58.37
0.08/99.67 13.72/87.12 0.01/99.79 4.76/99.66 0.09/99.97 3.18/97.83 3.25/99.48 1.78/92.82 2.20/98.70 5.04/98.46
SD v1.4 60.11 56.86 57.07 60.95 55.63 55.62 59.71 56.32 58.13 62.60
8.86/98.63 1.37/98.57 0.00/99.96 17.65/93.50 0.57/97.01 74.97/41.09 11.86/95.67 3.07/96.35 2.43/100.00 20.33/94.38
All GenImage 60.89 57.22 57.09 59.81 58.52 60.42 57.83 55.70 57.81 65.77
9.86/99.25 0.89/99.55 0.02/99.98 8.23/98.58 7.72/96.70 85.52/41.56 2.09/99.73 1.39/96.52 1.68/100.00 26.80/95.06
Table 6: Robustness on JPEG compression and Gaussian blur of AIDE. The accuracy (%) averaged over 16 test sets in $\mathcal{B}_{1}$ with specific perturbation.
Method Original JPEG Compression Gaussian Blur
QF=95 QF=90 QF=75 QF=50 $\sigma=1.0$ $\sigma=2.0$ $\sigma=3.0$ $\sigma=4.0$
CNNSpot 70.78 64.03 62.26 60.65 59.66 68.39 67.26 67.13 65.85
FreDect 64.03 66.95 67.45 66.64 65.33 65.75 66.48 68.58 69.64
Fusing 68.38 62.43 61.39 59.34 57.41 68.09 66.69 66.02 65.58
LNP 83.84 53.58 54.09 53.02 52.85 67.91 66.42 66.2 62.69
LGrad 75.34 51.55 51.39 50.00 50.00 71.73 69.12 68.43 66.22
DIRE-G 68.68 66.49 66.12 65.28 64.34 64.00 63.09 62.21 61.91
UnivFD 78.43 74.10 74.02 69.92 68.68 70.31 68.29 64.62 61.18
PatchCraft 89.31 72.48 71.41 69.43 67.78 75.99 74.90 73.53 72.28
AIDE 92.77 75.54 74.21 70.64 69.60 81.88 80.35 80.05 79.86
Table 7: Ablation studies on PFE and SFE modules.
@@
AIDE demonstrates a significant advancement over the existing state-of-the-art (SOTA) approach, for example, PatchCraft, achieving an average accuracy increase of 3.5%. UnivFD utilizes CLIP semantic features for detecting AI-generated images, proving effective for GAN-generated images. However, it shows pronounced performance degradation with diffusion model (DM)-generated images. This suggests that as generation quality improves, diffusion models produce images with fewer semantic artifacts, as depicted in Fig. 1 (a). Our approach, which integrates semantic, low-frequency, and high-frequency information at the feature level, enhances detection performance, yielding a 5.2% increase for GAN-based images and a 1.7% increase for DM-based images compared to the SOTA method.
On Benchmark Chameleon: As highlighted in Sec. 1, we contend that success on existing public benchmarks may not accurately reflect real-world scenarios or the advancement in AI-generated image detection, given that test sets are typically randomly sampled from generative models without “Turing Test”. To address potential biases related to training setups—such as generator types and image categories—we evaluate the performance of existing detectors under diverse training conditions. Despite their high performance on existing benchmarks, as depicted in Fig. 9, the state-of-the-art detector, PatchCraft, experiences substantial performance declines. Additionally, Table 5.2 reveals significant performance decreases across all methods, with most struggling to surpass an average accuracy close to random guessing (about 50%), indicating a failure in these contexts.
While our method achieves the SOTA results on available datasets, its performance on Chameleon remains low, especially on discovering the AI-generated images. This underscores that our dataset, Chameleon, which challenges human perception, is also extremely difficult AI models.
5.3 Robustness to Unseen Perturbations
In real-world scenarios, images inevitably encounter unseen perturbations in transmission and interaction, complicating AI-generated image detection. Herein, we assess the performance of various methods in handling potential perturbations, such as JPEG compression (Quality Factor (QF) = 95, 90, 75, 50) and Gaussian blur ($\sigma$ = 1.0, 2.0, 3.0, 4.0). As illustrated in Table 5.2, all methods exhibit a decline in performance due to disruptions in the pixel distribution. These disruptions diminish the discriminative artifacts left by generative models, complicating the differentiation between real and AI-generated images. Consequently, the robustness of these detectors in identifying AI-generated images is significantly compromised. Despite these challenging conditions, our method consistently outperforms others, maintaining a relatively higher accuracy in detecting AI-generated images. This superior performance is attributed to our model’s ability to effectively capture and leverage multi-perspective features, semantics, and noise, even when the pixel distribution is distorted.
Table 8: Performance of SOTA method, PatchCraft Zhong et al. (2023), under $\mathcal{B}_{1}$ (left), $\mathcal{B}_{2}$ (right), and our Chameleon testset. The boundary line for Acc = 50% is marked with a dashed line.
Table 9: Visualization of the effectiveness of PFE and SFE modules. Without SFE, AI-generated images exhibit pronounced semantic errors that are incorrectly classified as real. Similarly, without PFE, many AI-generated images, despite lacking semantic errors, contain subtle underlying noise that also leads to their misclassification as true.
5.4 Ablation Studies